OpenStudy (anonymous):

Solving the differential equation (2x+y+1)(dy/dx)=1

OpenStudy (anonymous):

I've thought about trying to write it as an exact equation, using an integrating factor. Of course it's not seperable. I can't see any way of writing it as a function f(x/y). I'm pretty shaking with this subject, so any help would be appreciated!

OpenStudy (anonymous):

IF YOU ARE GOOD AT ANALYZING POEMS GO TO http://www.twiddla.com/548645

OpenStudy (anonymous):

haha. No I'm not.

OpenStudy (anonymous):

is it the whole question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dy/dx = 1/2x+y+1 let P = 1/2x+y+1 = (2x+y+1)^-1

OpenStudy (anonymous):

So substitution for the whole expression on the right?

OpenStudy (anonymous):

find the dP/dx

OpenStudy (anonymous):

So isn't that the equivalent of taking the second derivative of y with respect to x?

OpenStudy (anonymous):

dP/dx in implicite, won't equivalent

OpenStudy (anonymous):

So now I have a second order differential equation \[d^2y/dx^2=dP/dx=-(2+ydy/dx)/(2x+y+1)\] ?

OpenStudy (anonymous):

Sorry the denominator on the right hand side of the expression should be squared

OpenStudy (anonymous):

in dP/dx result, dy/dx change into P and at this step we reach 1st order, and we shouldnt have y variable but in ur case why y cant be move out

OpenStudy (anonymous):

for example: P = 3x + 2y -10 dP/dx = 3+2 dy/dx F(p)= P = dy/dx dP/dx = 3 + 2 P integral (1/ (3+2P ) ) dP = integral (dx)

OpenStudy (anonymous):

sorry - how does F(p)=P?

OpenStudy (anonymous):

Otherwise...yeah I see what's going on. Its a good trick. But I guess I have to find a different substitution to solve this equation?

OpenStudy (anonymous):

sorry because im not english native, so hard for me to explain

OpenStudy (anonymous):

hey, that's ok. I really do appreciate that you're trying!

OpenStudy (anonymous):

check this out http://www.wolframalpha.com/input/?i=%282x%2By%2B1%29%28dy%2Fdx%29%3D1